# Bangle functions are the generic basis for cluster algebras from punctured surfaces with boundary

### v2

@@ 1.1. Bases for cluster algebras
Since the set of cluster monomials is always linearly independent [4] this led more generally to the search of bases for (upper) cluster algebras which contain the cluster monomials, and which ideally expand positively with respect to each cluster. Moreover, such ”good” bases are typically parametrized by the by tropical points of the corresponding cluster Poisson variety. By the landmark work of Gross–Hacking–Keel–Kontsevich [28] on theta functions at least one such basis exists under rather weak hypothesis. See the remarkable paper [49] for a thorough discussion.
@@
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a (bordered) surface with marked points in the sense of [15], where $\mathbb{M}$ denotes the set of marked points on the boundary and $\mathbb{P}$ is the set of punctures. See Definition 2.5 for more details. To associate a (coefficient-free) cluster algebra to $\mathbf{\Sigma}$ one considers tagged triangulations of $\mathbf{\Sigma}$. Following [15, Def. 9.6] each tagged triangulation $T$ has an associated skew-symmetric matrix $B(T)$, see also Definition 2.9 for details. The matrix $B(T)$ changes under any flip of an arc in $T$ according to the Fomin–Zelevinsky mutation rule. This allows us already to define a cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{A}(B(T),\mathbf{x})$ which is actually independent of the choice of the triangulation $T$. In fact, a fundamental result of Fomin–Shapiro–Thurston [16, Cor. 6.2] and [15, Thm. 7.11] states that provided $\mathbf{\Sigma}$ is not a closed surface with exactly one puncture, there is a bijection $\gamma\mapsto x_{\gamma}$ between the set of tagged arcs on $\mathbf{\Sigma}$ and the set of cluster variables of $\mathcal{A}(B(T_{0}),\mathbf{x})$, such that the induced assignment $T\mapsto(B(T),(x_{\gamma}))_{\gamma_{\in}T}$ is a bijection between the set of triangulations of $\mathbf{\Sigma}$ and the set of (unlabelled) seeds of $\mathcal{A}(B(T_{0}),\mathbf{x})$, such that whenever two tagged triangulations are related by the flip of a tagged arc, the seeds corresponding to them are related by the mutation of seeds corresponding to the arc being flipped.
Given a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ and a tagged triangulation $T$ of $\mathbf{\Sigma}$, Musiker–Schiffler–Williams associated to each tagged arc (resp. simple closed curve) $\alpha$ on $\mathbf{\Sigma}$, a bipartite graph $G(T,\alpha)$, which they called snake graph (resp. band graph) [43, 44], with edges and tiles naturally labelled by the arcs in $T$ (and the boundary segments of $\mathbf{\Sigma}$). Using these labels, they assigned a weight monomial $w(P):=x(P)y(P)$ to each perfect (resp. good) matching $P$ of $G(T,\alpha)$, thus producing a generating function $\sum_{P}w(P)$ for the perfect matchings (resp. good matchings) of $G(T,\alpha)$. Dividing this generating function by the monomial $\operatorname{cross}(T,\alpha)$ that records the number of crossings of $\alpha$ with each arc in $T$, they define the bangle function
@@
Let $Q$ be a 2-acyclic quiver with vertex set $Q_{0}=\{1,2,\ldots,n\}$, and $B(Q)=(b_{ij})\in\mathbb{Z}^{Q_{0}\times Q_{0}}$ the corresponding skew-symmetric matrix with with entries
@@
as in [10, Equation (1.4)]. In this context we abbreviate $\mathcal{A}(Q):=\mathcal{A}(B(Q))$ for the corresponding cluster algebra with trivial coefficients.
@@
Together with Leclerc and Schröer, the first named author explored in [25] the algebraic geometry of (affine) varieties of representations of Jacobi-finite non-degenerate quivers with potential $(Q,S)$ which appear in algebraic Lie theory. They isolated a class of irreducible components of these representation spaces with good geometric-homological properties and called them strongly reduced irreducible components. Later on, the adjective used was changed to generically $\tau^{-}$-reduced, or simply, $\tau^{-}$-reduced, cf. [23]. Roughly speaking, a component $Z\subseteq\operatorname{rep}(\mathcal{P}(Q,S),\mathbf{d})$ is $\tau^{-}$-reduced if the codimension in $Z$ of a top dimensional $\operatorname{GL}_{\mathbf{d}}(\mathbb{C})$-orbit $\mathcal{O}\subseteq Z$ is equal to the minimum value taken on $Z$ by Derksen-Weyman-Zelevinsky’s (injective) $E$-invariant. See [6, Sec. 3 and 5] for more details.
It is easy to see that each $\tau^{-}$-reduced irreducible component $Z$ has an open dense subset where the constructible function $M\mapsto CC_{\mathcal{P}(Q,S)}(M)$ takes a constant value, which is then denoted $CC_{\mathcal{P}(Q,S)}(Z)$. Resorting to decorated representations and spaces of decorated representations in order to be able to hit initial cluster variables through the Caldero-Chapoton function, the set of generic Caldero-Chapoton functions is
[…]
@@ 1.4. Statement of the Main Result
Let $\mathbf{\Sigma}$ be a surface with marked points. To each tagged triangulation $T$ of $\mathbf{\Sigma}$ we associate a quiver $Q^{\prime}(T)$ such that $B(Q^{\prime}(T))=B(T)$. This convention allows us to state and prove our results without the roundabout with dual CC-functions as for example in [23]. Schröer and the first two named authors showed [22, Thm. 1.18], relying heavily on results of Mills [40], Muller [41, 42] and Qin [49], that when $(Q,S)=(Q^{\prime}(T),S^{\prime}(T))$ for some triangulation $T$ of a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ with at least two marked points on the boundary, the set of generic CC-functions $\mathcal{B}_{\mathcal{P}(Q,S)}$ is indeed a basis for the (upper) cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$, where $S^{\prime}(T)$ is the potential defined by Cerulli Irelli and the second author in [5]. Let us write from now on $A(T)$ for the corresponding complex Jacobian algebra $\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T),W^{\prime}(T))$. Note that due to our convention here, the algebra $A(T)$ is opposite to the one which was considered for example in [24]. We are now ready to state our main result, Theorem 6.3, in a slightly simplified version.
@@
(1) The proof of our theorem relies heavily on [24] and our Combinatorial Key Lemma 5.3. These arguments imply $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$. See Section 1.5 for more details. By the above mentioned result from  [22, Thm. 1.18], the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ is indeed a basis of the upper cluster algebra.
@@
(3) Schröer and the first two named authors showed in [23, Thm. 11.9] this theorem for the special case $\mathbb{P}=\emptyset$. Their proof relies heaavily on the fact that for $\mathbb{P}=\emptyset$ the Jacobian algebra $A(T)$ is gentle, which allows for quite direct calculations. In our more general situation the algebras $A(T)$ are still tame [21], but the description of their indecomposable representations is in general unknown. See for example [11] for partial results into this direction. Our approach here exploits the use of mutations in the spirit of [10].
@@ 1.5. Structure of the paper
The paper is organized as follows. In Section 2 we provide the basics about mutations and cluster algebras from surfaces. In Section 3 we summarize the main constructions and results from [24] upon which we will rely in Section 6, the most relevant one being [24, Proposition 2.8] (stated as Proposition 3.8 below), which gives a precise recursive formula for the change that the generic projective $g$-vector of any $\tau$-reduced component undergoes under mutation of $\tau$-reduced components.
It should be stressed that, since elements of $\mathcal{B}^{\circ}$ were defined directly from the associated snake or band graphs, then, a priori, this viewpoint says nothing about whether $\mathcal{B}^{\circ}$ is canonical in a cluster-theoetrical sense. That is, whether or not $\mathcal{B}^{\circ}$ is independent of the choice of triangulation, up to a cluster isomorphism.
@@
(1.5) $g_{j}^{\prime}:=\left\{\begin{array}[]{ll}-g_{k},&\text{if $j=k$};\\ \\ g_{j}+[b_{jk}]_{+}g_{k}-b_{jk}h_{k},&\text{if $j\neq k$}.\\ \end{array}\right.$
@@
  • the orbit closure of $\mathcal{M}(T_{\alpha},\alpha)$ is a $\tau$-reduced component $Z_{T_{\alpha},\alpha}$ and any decorated representation mutation-equivalent to $\mathcal{M}(T_{\alpha},\alpha)$ is generic, for the values of the projective $g$-vector, the representation-theoretic (dual) $F$-polynomial and Derksen-Weyman-Zelevinsky’s $E$-invariant, in the corresponding $\tau$-reduced component.
@@ 2.1. Cluster algebras
  • $\mathbf{x}=(x_{1},\ldots x_{n})$ is a collection of variables in $\mathcal{F}$ which are algebraically independent over $\mathbb{ZP}$,
@@
  • $\mathbf{x}^{\prime}=(x^{\prime}_{1},\ldots x^{\prime}_{n})$ is defined by
@@
    $b^{\prime}_{jk}=\begin{cases}-b_{jk},&\text{if $j=i$ or $k=i$,}\\ b_{jk}+\max(0,-b_{ji})b_{ik}+\max(0,b_{ik})b_{ji},&\text{otherwise.}\\ \end{cases}$
@@
Given a skew-symmetric $n\times n$ matrix $B$ we define a quiver $Q(B)=(Q_{0},Q_{1})$ by setting $Q_{0}=\{1,\ldots,n\}$ and demanding that for any $i,j\in Q_{0}$ there are $[b_{ij}]_{+}$ arrows from $j$ to $i$ in $Q_{1}$.
@@ 2.2. Cluster algebras from surfaces
A surface with marked points is a triple $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ consisting of a compact, connected, oriented two-dimensional real differentiable manifold $\Sigma$ with (possibly empty) boundary $\partial\Sigma$, a finite set $\mathbb{M}\subseteq\partial\Sigma$ containing at least one point from each boundary component of $\Sigma$, and a finite set $\mathbb{P}\subseteq\Sigma\setminus\partial\Sigma$. It is required that $\mathbb{M}\cup\mathbb{P}\neq\varnothing$. The elements of $\mathbb{M}\cup\mathbb{P}$ are called marked points and the elements of $\mathbb{P}$ are called punctures. If $\mathbb{P}=\varnothing$, it is said that $\mathbf{\Sigma}$ is unpunctured, whereas if $\mathbb{P}\neq\varnothing$, one says that $\mathbf{\Sigma}$ is punctured. For technical reasons we exclude the cases where $\mathbf{\Sigma}$ is an unpunctured or once-punctured monogon; a digon; a triangle; or a once, twice or thrice punctured sphere. In [15, Def. 2.1] the term bordered surfaces with marked points was used instead of the more recent terminology surface with marked points or even marked surface.
@@
$\pi_{T}(i)=\begin{cases}j&\text{if $i$ is the folded side of a self-folded triangle in $T^{\circ}$,}\\ &\text{and $j$ is the remaining side;}\\ i&\text{otherwise.}\\ \end{cases}$
@@
$b^{\Delta}_{jk}=\begin{cases}1&\text{if $\Delta$ has sides $\pi_{T}(j)$ and $\pi_{T}(k)$, and $\pi_{T}(k)$ follows $\pi_{T}(j)$}\\ &\text{in the clockwise sense;}\\ -1&\text{if $\Delta$ has sides $\pi_{T}(j)$ and $\pi_{T}(k)$, and $\pi_{T}(k)$precedes $\pi_{T}(j)$}\\ &\text{in the clockwise sense;}\\ 0&\text{otherwise.}\\ \end{cases}$
@@
[…]
@@
3. Laminations as $\tau$-reduced components
@@ 3.1. Tame partial KRS-monoids
The simple-minded notion of partial Krull-Remak-Schmidt monoid was introduced in [24] as a convenient abstract framework to state the naturalness of the bijection between laminations and $\tau$-reduced components proved therein.
@@
[24, Definition 2.1] A partial monoid is a triple $(X,e,\oplus)$ consisting of a set $X$, a symmetric function $e:X\times X\rightarrow\mathbb{N}$ and a partially-defined sum :$\oplus:\{(x,y)\in X\times X\mid e(x,y)=0\}\rightarrow X$ such that:
@@
[24, Definition 2.3] Suppose $X=(X,e,\oplus)$ is a partial monoid.
@@
[24, Example 2.4] Suppose $C$ is a set equipped and $e:C\times C\rightarrow\mathbb{N}$ is a function such that $e(c,c)=0$ for all $c\in C$. Then
$\operatorname{KRS}(C,e):=\{f:C\rightarrow\mathbb{N}\mid c_{1},c_{2}\in\operatorname{supp}(f)\Rightarrow e(c_{1},c_{2})=0\text{ and }\lvert\operatorname{supp}(f)\rvert<\infty\}$
@@
3.2. KRS-monoids of $\tau^{-}$-reduced components
Let $A$ be a basic finite dimensional algebra over an algebraically closed field $𝕜\Bbbk$. We may identify the Grothendieck group $K_{0}(A)$ of $A$ with $\mathbb{Z}^{n}$ for some non-negative integer $n$. Consider the set $\operatorname{DecIrr}^{\tau^{-}}(A)$ of decorated, generically $\tau^{-}$-reduced, irreducible components of the representation varieties of $A$. See Section 1.3 for the terminology, and [6, Sec. 3 and 5] for more details. The set $\operatorname{DecIrr}^{\tau^{-}}(A)$ becomes a partial KRS-monoid under the function

### v3

@@ 1.1. Bases for cluster algebras
Since the set of cluster monomials is always linearly independent [5] this led more generally to the search of bases for (upper) cluster algebras which contain the cluster monomials, and which ideally expand positively with respect to each cluster. Moreover, such “good” bases are typically parametrized by the by tropical points of the corresponding cluster Poisson variety. By the landmark work of Gross–Hacking–Keel–Kontsevich [29] on theta functions at least one such basis exists under rather weak hypothesis. See the remarkable paper [50] for a thorough discussion.
@@
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a (bordered) surface with marked points in the sense of [16], where $\mathbb{M}$ denotes the set of marked points on the boundary and $\mathbb{P}$ is the set of punctures. See Definition 2.5 for more details. To associate a (coefficient-free) cluster algebra to $\mathbf{\Sigma}$ one considers tagged triangulations of $\mathbf{\Sigma}$. Following [16, Def. 9.6] each tagged triangulation $T$ has an associated skew-symmetric matrix $B(T)$, see also Definition 2.9 for details. The matrix $B(T)$ changes under any flip of an arc in $T$ according to the Fomin–Zelevinsky mutation rule. This allows us already to define a cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{A}(B(T),\mathbf{x})$ which is actually independent of the choice of the triangulation $T$. In fact, a fundamental result of Fomin–Shapiro–Thurston [17, Cor. 6.2] and [16, Thm. 7.11] states that provided $\mathbf{\Sigma}$ is not a closed surface with exactly one puncture, there is a bijection $\gamma\mapsto x_{\gamma}$ between the set of tagged arcs on $\mathbf{\Sigma}$ and the set of cluster variables of $\mathcal{A}(B(T_{0}),\mathbf{x})$, such that the induced assignment $T\mapsto(B(T),(x_{\gamma}))_{\gamma_{\in}T}$ is a bijection between the set of triangulations of $\mathbf{\Sigma}$ and the set of (unlabelled) seeds of $\mathcal{A}(B(T_{0}),\mathbf{x})$, such that whenever two tagged triangulations are related by the flip of a tagged arc, the seeds corresponding to them are related by the mutation of seeds corresponding to the arc being flipped.
Given a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ and an ideal triangulation $T$ of $\mathbf{\Sigma}$, Musiker–Schiffler–Williams associated to each plain arc (resp. simple closed curve) $\alpha$ on $\mathbf{\Sigma}$, a bipartite graph $G(T,\alpha)$, which they called a snake graph (resp. band graph) [44, 45], with edges and tiles naturally labelled by the arcs in $T$ (and the boundary segments of $\mathbf{\Sigma}$). Later, Wilson extended this construction to all tagged arcs $\alpha$, with respect to any tagged triangulation $T$, introducing the like-minded loop graph $G(T,\alpha)$ [52]. Considering a certain subset $\mathcal{P}(G(T,\alpha))$ of the set of all perfect matchings of the loop or band graph $G(T,\alpha)$, called good matchings, and using the natural labels mentioned above, they assigned a weight monomial $w(P):=x(P)y(P)$ to each $P\in\mathcal{P}(G(T,\alpha))$, thus producing a generating function $\sum_{P}w(P)$ for the good matchings of $G(T,\alpha)$. Dividing this generating function by the monomial $\operatorname{cross}(T,\alpha)$ that records the number of crossings of $\alpha$ with each arc in $T$, they define the bangle function
@@
Let $Q$ be a 2-acyclic quiver with vertex set $Q_{0}=\{1,2,\ldots,n\}$, and $B(Q)=(b_{ij})\in\mathbb{Z}^{Q_{0}\times Q_{0}}$ the corresponding skew-symmetric matrix with entries
@@
as in [11, Equation (1.4)]. In this context we abbreviate $\mathcal{A}(Q):=\mathcal{A}(B(Q))$ for the corresponding cluster algebra with trivial coefficients.
@@
Together with Leclerc and Schröer, the first named author explored in [26] the algebraic geometry of (affine) varieties of representations of Jacobi-finite non-degenerate quivers with potential $(Q,S)$ which appear in algebraic Lie theory. They isolated a class of irreducible components of these representation spaces with good geometric-homological properties and called them strongly reduced irreducible components. Later on, the adjective used was changed to generically $\tau^{-}$-reduced, or simply, $\tau^{-}$-reduced, cf. [24]. Schröer and Bobiński made in a recent preprint [1] a very good point that these components should in fact be called $\tau^{-}$-regular components. We decided to use this new terminology. Roughly speaking, a component $Z\subseteq\operatorname{rep}(\mathcal{P}(Q,S),\mathbf{d})$ is $\tau^{-}$-regular if the codimension in $Z$ of a top dimensional $\operatorname{GL}_{\mathbf{d}}(\mathbb{C})$-orbit $\mathcal{O}\subseteq Z$ is equal to the minimum value taken on $Z$ by Derksen-Weyman-Zelevinsky’s (injective) $E$-invariant. See [7, Sec. 3 and 5] for more details.
It is easy to see that each $\tau^{-}$-regular irreducible component $Z$ has an open dense subset where the constructible function $M\mapsto CC_{\mathcal{P}(Q,S)}(M)$ takes a constant value, which is then denoted $CC_{\mathcal{P}(Q,S)}(Z)$. Resorting to decorated representations and spaces of decorated representations in order to be able to hit initial cluster variables through the Caldero-Chapoton function, the set of generic Caldero-Chapoton functions is
$\mathcal{B}_{\mathcal{P}(Q,S)}:=\{CC_{\mathcal{P}(Q,S)}(Z)\ |\ \text{$Z$ is a decorated $\tau^{-}$-regular irreducible component}\}.$
@@ 1.4. Statement of the Main Result
Let $\mathbf{\Sigma}$ be a surface with marked points. To each tagged triangulation $T$ of $\mathbf{\Sigma}$ we associate a quiver $Q^{\prime}(T)$ such that $B(Q^{\prime}(T))=B(T)$. This convention allows us to state and prove our results without the roundabout with dual CC-functions as for example in [24]. Schröer and the first two named authors showed [23, Thm. 1.18], relying heavily on results of Mills [41], Muller [42, 43] and Qin [50], that when $(Q,S)=(Q^{\prime}(T),S^{\prime}(T))$ for some triangulation $T$ of a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ with at least two marked points on the boundary, the set of generic CC-functions $\mathcal{B}_{\mathcal{P}(Q,S)}$ is indeed a basis for the (upper) cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$, where $S^{\prime}(T)$ is the potential defined by Cerulli Irelli and the second author in [6]. Let us write from now on $A(T)$ for the corresponding complex Jacobian algebra $\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T),W^{\prime}(T))$. Note that due to our convention here, the algebra $A(T)$ is opposite to the one which was considered for example in [25]. We are now ready to state our main result, Theorem 6.3, in a slightly simplified version.
@@
(1) The proof of our theorem relies heavily on [25] and our Combinatorial Key Lemma 5.3. These arguments imply $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$. See Section 1.5 for more details. By the above mentioned result from  [23, Thm. 1.18], the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ is indeed a basis of the upper cluster algebra.
@@
(3) Schröer and the first two named authors showed in [24, Thm. 11.9] this theorem for the special case $\mathbb{P}=\emptyset$. Their proof relies heavily on the fact that for $\mathbb{P}=\emptyset$ the Jacobian algebra $A(T)$ is gentle, which allows for quite direct calculations. In our more general situation the algebras $A(T)$ are still tame [22], but the description of their indecomposable representations is in general unknown. See for example [12] for partial results into this direction. Our approach here exploits the use of mutations in the spirit of [11].
@@ 1.5. Structure of the paper
The paper is organized as follows. In Section 2 we provide the basics about mutations and cluster algebras from surfaces. In Section 3 we summarize the main constructions and results from [25] upon which we will rely in Section 6, the most relevant one being [25, Proposition 2.8] (stated as Proposition 3.8 below), which gives a precise recursive formula for the change that the generic projective $g$-vector of any $\tau$-regular component undergoes under mutation of $\tau$-regular components.
It should be stressed that, since elements of $\mathcal{B}^{\circ}$ were defined directly from the associated loop or band graphs, then, a priori, this viewpoint says nothing about whether $\mathcal{B}^{\circ}$ is canonical in a cluster-theoretical sense. That is, whether or not $\mathcal{B}^{\circ}$ is independent of the choice of triangulation, up to a cluster isomorphism.
@@
[…]
@@
  • the orbit closure of $\mathcal{M}(T_{\alpha},\alpha)$ is a $\tau$-regular component $Z_{T_{\alpha},\alpha}$ and any decorated representation mutation-equivalent to $\mathcal{M}(T_{\alpha},\alpha)$ is generic, for the values of the projective $g$-vector, the representation-theoretic (dual) $F$-polynomial and Derksen-Weyman-Zelevinsky’s $E$-invariant, in the corresponding $\tau$-regular component.
@@ 2.1. Cluster algebras
  • $\mathbf{x}=(x_{1},\ldots,x_{n})$ is a collection of variables in $\mathcal{F}$ which are algebraically independent over $\mathbb{ZP}$,
@@
  • $\mathbf{x}^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{n})$ is defined by
@@
[…]
@@
Given a skew-symmetric $n\times n$ matrix $B$ we define a quiver $Q(B)=(Q_{0},Q_{1})$ by setting $Q_{0}=\{1,\ldots,n\}$ and demanding that for any $i,j\in Q_{0}$ there are $b_{ij}]_{+}$ arrows from $j$ to $i$ in $Q_{1}$.
@@ 2.2. Cluster algebras from surfaces
A surface with marked points is a triple $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ consisting of a compact, connected, oriented two-dimensional real differentiable manifold $\Sigma$ with (possibly empty) boundary $\partial\Sigma$, a finite set $\mathbb{M}\subseteq\partial\Sigma$ containing at least one point from each boundary component of $\Sigma$, and a finite set $\mathbb{P}\subseteq\Sigma\setminus\partial\Sigma$. It is required that $\mathbb{M}\cup\mathbb{P}\neq\varnothing$. The elements of $\mathbb{M}\cup\mathbb{P}$ are called marked points and the elements of $\mathbb{P}$ are called punctures. If $\mathbb{P}=\varnothing$, it is said that $\mathbf{\Sigma}$ is unpunctured, whereas if $\mathbb{P}\neq\varnothing$, one says that $\mathbf{\Sigma}$ is punctured. For technical reasons we exclude the cases where $\mathbf{\Sigma}$ is an unpunctured or once-punctured monogon; a digon; a triangle; or a once, twice or thrice punctured sphere. In [16, Def. 2.1] the term bordered surfaces with marked points was used instead of the more recent terminology surface with marked points or even marked surface.
@@
[…]
@@
[…]
@@
Cluster variables $\longleftrightarrow$ Tagged arcs
Clusters $\longleftrightarrow$ Tagged triangulations
Mutation $\longleftrightarrow$ Flips of tagged arcs
@@
3. Laminations as $\tau$-regular components
@@ 3.1. Tame partial KRS-monoids
The simple-minded notion of partial Krull-Remak-Schmidt monoid was introduced in [25] as a convenient abstract framework to state the naturalness of the bijection between laminations and $\tau$-regular components proved therein.
@@
[25, Definition 2.1] A partial monoid is a triple $(X,e,\oplus)$ consisting of a set $X$, a symmetric function $e:X\times X\rightarrow\mathbb{N}$ and a partially-defined sum $\oplus:\{(x,y)\in X\times X\mid e(x,y)=0\}\rightarrow X$ such that:
@@
[25, Definition 2.3] Suppose $X=(X,e,\oplus)$ is a partial monoid.
@@
[25, Example 2.4] Suppose $C$ is a set equipped and $e:C\times C\rightarrow\mathbb{N}$ is a function such that $e(c,c)=0$ for all $c\in C$. Then
$\operatorname{KRS}(C,e):=\{f:C\rightarrow\mathbb{N}\mid c_{1},c_{2}\in\operatorname{supp}(f)\Rightarrow e(c_{1},c_{2})=0\text{ and }\lvert\operatorname{supp}(f)\rvert$$\infty\}$
@@
3.2. KRS-monoids of $\tau^{-}$-regular components
[…]